# Meta-Learning Objectives for Preference Optimization

### v2

@@ Abstract
Evaluating preference optimization (PO) algorithms on LLM alignment is a challenging task that presents prohibitive costs, noise, and several variables like model size and hyper-parameters. In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on much simpler benchmarks. We design a diagnostic suite of MuJoCo tasks and datasets, which we use to systematically evaluate PO algorithms, establishing a more controlled and cheaper benchmark. We then propose a novel family of PO algorithms based on mirror descent, which we call Mirror Preference Optimization (MPO). Through evolutionary strategies, we search this class to discover algorithms specialized to specific properties of preference datasets, such as mixed-quality or noisy data. We demonstrate that our discovered PO algorithms outperform all known algorithms in the targeted MuJoCo settings. Finally, based on the insights gained from our MuJoCo experiments, we design a novel PO algorithm that significantly outperforms existing baselines in an LLM alignment task.
@@ 1 Introduction
For each setting we consider, we discover an algorithm that significantly outperforms all baselines. Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one. We use this insight to design a new PO function, Constrained Linear Policy Optimization (CoLPO) that demonstrates promising results in an LLM alingment task. We summarize our contributions below.
@@
    We perform a systematic evaluation of eight of the most popular PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy. We see that most existing algorithms struggle when dealing with noise and mixed-quality date.
@@
    We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks. In particular, we show that our CoLPO algorithm significantly improves upon the ORPO, CPO, and SLiC-HF baselines.
Overall, we show that it is possible to gain relevant insights for the design of LLM alignment algorithms through a systematic analysis in simple, computationally inexpensive environments such as MuJoCo.
@@ 2 Preliminaries
A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$. Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$. Lastly, let $\pi(\cdot|\tau)$ be a distribution over $(\Delta(\mathcal{A}))^{T}$ defined as $\pi(\cdot|s_{0})\times\dots\times\pi(\cdot|s_{N-1})$.
@@
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)$ (1)
$:=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
where $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$. Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$. Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}$ (2)
$=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$
where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter. In this work, we consider an offline training setting, where the agent aim to solve the optimization problem in (1) but only has access to the the dataset $\mathcal{D}$ and cannot collect further data. We also assume the agent does not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
@@ 2.1 Alignment to preference feedback
SFT is an initial alignment phase, where the policy $\pi_{0}$ is trained to imitate high-quality demonstration data. In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$. We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure.
@@
$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}g\circ\sigma((r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l}))/\eta),$ (3)
for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$, for an increasing function $g:\mathbb{R}\leftrightarrow\mathbb{R}$, and for $\eta\geq 0$. Here, typically, $g$ is set to be the logarithm function. The policy optimization problem is the one to maximize the expected reward, that is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (4)
$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$
@@
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$ (5)
$\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$
which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3), with $g=\log$. Refer to Appendix C for details. Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs (Yuan et al., 2024; Jiang et al., 2024).
@@
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg{[}\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}$ (6)
$+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg{]}$
@@
Research on preference optimization has been very active and many methods have been proposed. We present a summary of the most popular algorithms in Table 1 and a brief discussion in Appendix A.
Beyond these implicit algorithms, there are several other methods that explicitly solve the maximum likelihood problem in (3) and use the learned reward model to optimize the objective in (4) with an RL algorithm. Overall, RLHF is a superior approach and the industry standard, but is more computationally expensive and complex to implement. For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix F.
Method Objective
IPO (Azar et al., 2024) $-\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
Table 1: A summary of popular PO algorithms. The objective in each row is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.
@@ 2.2 Mirror Maps
We review the concept of mirror map, which will be needed when describing our methodology. For a convex set $\mathcal{X}\subseteq\mathbb{R}^{|\mathcal{A}|}$, a mirror map $h:\mathcal{X}\rightarrow\mathbb{R}$ is defined as a strictly convex, continuously differentiable and essentially smooth function [^1] A function $h$ is essentially smooth if $\lim_{x\rightarrow\partial\mathcal{X}}\lVert\nabla h(x)\rVert_{2}=+\infty$, where $\partial\mathcal{X}$ denotes the boundary of $\mathcal{X}$. function that satisfies $\nabla h(\mathcal{X})=\mathbb{R}^{|\mathcal{A}|}$. Essentially, a mirror map is a function whose gradient allows bijective mapping between the primal space $\mathcal{X}$ and the dual space $\mathbb{R}^{|\mathcal{A}|}$. The specific class of mirror maps that we are going to use is the $\omega$-potential mirror map class, to which most mirror maps considered in the literature belong.
Definition 2.1 ($\omega$-potential mirror map (Krichene et al., 2015)).
@@
[…]
For any $\omega$-potential $\phi$, we define the associated mirror map $h_{\phi}$ as
$h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix D). Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one. Additionally, the same $\omega$-potential $\phi$ can be used to generate mirror maps for different action spaces, allowing the insights obtained for one action space to easily generalize to others. An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))$
$-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$. When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix D). When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
@@ 3 Mirror Preference Optimization
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (7)
$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$
@@
Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
@@
We provide a proof for Theorem 3.1 in Appendix C. By plugging (11) in the maximum likelihood problem in (3), we obtain the objective:
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}g\circ\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))$ (9)
$-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$
@@
Two-step MPO
We can use the objective in (9) to define a class of two-step PO algorithms, which consist of a preliminary SFT phase to obtain the reference policy $\pi_{\mathrm{ref}}$ and a PO phase which optimizes (9). When $\phi=e^{x}$, the (9) is equivalent to (5) and we recover the DPO algorithm.
One-step MPO
@@
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\psi(\pi(\tau_{w}))$ (10)
$+\lambda g\circ\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big{]},$
where $\psi$ is an $\omega$-potential. Here, the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ have canceled out due to $\pi_{\mathrm{ref}}$ being uniform. We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization. When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (10) recovers the ORPO objective in (6).
The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification. In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform current baselines.

### v3

@@ Abstract
Evaluating preference optimization (PO) algorithms on LLM alignment is a challenging task that presents prohibitive costs, noise, and several variables like model size and hyper-parameters. In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on simpler benchmarks. We design a diagnostic suite of MuJoCo tasks and datasets, which we use to systematically evaluate PO algorithms, establishing a more controlled and cheaper benchmark. We then propose a novel family of PO algorithms based on mirror descent, which we call Mirror Preference Optimization (MPO). Through evolutionary strategies, we search this class to discover algorithms specialized to specific properties of preference datasets, such as mixed-quality or noisy data. We demonstrate that our discovered PO algorithms outperform all known algorithms in the targeted MuJoCo settings. Finally, based on the insights gained from our MuJoCo experiments, we design a PO algorithm that significantly outperform existing baselines in an LLM alignment task.
11footnotetext: Equal contribution, order decided by coin flip.22footnotetext: Now at Amazon.
@@ 1 Introduction
For each setting we consider, we discover an algorithm that significantly outperforms all baselines. Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one. We use this insight to design a new PO algorithm, Temporally-Aware Mirror Preference Optimization (TA-MPO), which demonstrate promising results in an LLM alignment task. We summarize our contributions below.
@@
    We perform a systematic evaluation of eight existing PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy. We see that most existing algorithms struggle when dealing with noise and mixed-quality data.
@@
    We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks. In particular, we show that our TA-MPO algorithm significantly improves upon the baselines.
@@ 2 Preliminaries
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,T,\mu)$ denote an episodic Markov Decision Process, where $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action spaces, $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ when taking action $a$, $r(s,a)\in[0,1]$ is the reward function, $T$ is the maximum episode length, and $\mu$ is a starting state distribution. A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$. Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$. Lastly, let $\pi(\cdot|\tau)$ be a distribution over $(\Delta(\mathcal{A}))^{T}$ defined as $\pi(\cdot|s_{0})\times\dots\times\pi(\cdot|s_{N-1})$.
@@
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$ (1)
where $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$. Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$. Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w}))}{\exp(r(\tau_{w}))+\exp(r(\tau_{l}))}=\sigma(r(\tau_{w})-r(\tau_{l})),$ (2)
where $\sigma$ is the sigmoid function. In this work, we consider an offline training setting, where the agent aim to solve the optimization problem in (1) but only has access to the the dataset $\mathcal{D}$ and cannot collect further data. We also assume the agent does not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
@@ 2.1 Alignment to preference feedback
SFT is an initial alignment phase, where the policy $\pi_{0}$ is trained to imitate high-quality demonstration data. The starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$. We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure.
@@
$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\sigma(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})),$ (3)
for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$. The policy optimization problem is the one to maximize the expected reward, that is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$ (4)
@@
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$ (5)
which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3). Refer to Appendix˜C for details. Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs (Yuan et al., 2024; Jiang et al., 2024).
@@
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg[\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg]$ (6)
@@
Research on preference optimization has been very active and many methods have been proposed. We present a summary of some among the most popular algorithms in Table˜4 and a brief discussion in Appendix A. Beyond these implicit algorithms, there are several other methods that explicitly solve the maximum likelihood problem in (3) and use the learned reward model to optimize the objective in (4) with an RL algorithm. Overall, RLHF is a superior approach and the industry standard, but is more computationally expensive and complex to implement. For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix H.
@@ 2.2 Mirror Maps
We review the concept of mirror map, which will be needed when describing our methodology. For a convex set $\mathcal{X}\subseteq\mathbb{R}^{|\mathcal{A}|}$, a mirror map $h:\mathcal{X}\rightarrow\mathbb{R}$ is defined as a strictly convex, continuously differentiable and essentially smooth function***A function $h$ is essentially smooth if $\lim_{x\rightarrow\partial\mathcal{X}}\lVert\nabla h(x)\rVert_{2}=+\infty$, where $\partial\mathcal{X}$ denotes the boundary of $\mathcal{X}$. function that satisfies $\nabla h(\mathcal{X})=\mathbb{R}^{|\mathcal{A}|}$. Essentially, a mirror map is a function whose gradient allows bijective mapping between the primal space $\mathcal{X}$ and the dual space $\mathbb{R}^{|\mathcal{A}|}$. The specific class of mirror maps that we are going to use is the $\omega$-potential mirror map class, to which most mirror maps considered in the literature belong.
Definition 2.1 ($\omega$-potential mirror map Krichene et al. (2015)).
@@
$\lim_{x\rightarrow-\infty}\phi(x)=\omega,\;\lim_{x\rightarrow u}\phi(x)=+\infty,\;\int_{0}^{1}\phi^{-1}(x)dx\leq\infty.$
For any $\omega$-potential $\phi$, the associated mirror map is $h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix˜F). Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one. Additionally, the same $\omega$-potential $\phi$ can be used to generate mirror maps for different action spaces, allowing the insights obtained for one action space to easily generalize to others. An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$. When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix˜F). When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
@@ 3 Mirror Preference Optimization
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\bigg[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\bigg],$ (7)
@@
Let $h_{\phi}$ be a 0-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
$r(\tau)=\beta\phi^{-1}(\pi^{\star}(\tau))-\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (8)
@@
We provide a proof for Theorem˜3.1 in Appendix˜C. The next step is to model the reward using a classification problem based on the reward difference rather than the maximum likelihood problem in (3), as suggested by Tang et al. (2024). That is, our aim is to solve the optimization problem
$\widehat{r}\in\mathop{\mathrm{argmax}}\nolimits_{r_{\theta}}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}g(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})),$ (9)
where $g$ is an increasing function. We give further details on this interpretation of reward modeling in Appendix˜D. By plugging (8) in the optimization problem in (9), we obtain the objective:
$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{\mathcal{D}}\big[g\big(\beta(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big)\big],$ (10)
@@
Two-step MPO (2S-MPO)
We can use the objective in (10) to define a class of two-step PO algorithms, which consist of a preliminary SFT phase to obtain the reference policy $\pi_{\mathrm{ref}}$ and a PO phase which optimizes (10). When $\phi=e^{x}$, the (10) is equivalent to (5) and we recover DPO.
One-step MPO (1S-MPO)
@@
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[\psi(\pi(\tau_{w}))+\lambda g\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big],$ (11)
where $\psi$ is an $\omega$-potential. The term $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ has canceled out due to $\pi_{\mathrm{ref}}$ being uniform. We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization. When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (11) recovers the ORPO objective in (6).
Temporally-Aware MPO (TA-MPO)
Lastly we design a variation of MPO that gradually switches from SFT to PO. TA-MPO consists of single-phase algorithms that optimize the objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[(1-\alpha(t))\psi(\pi(\tau_{w}))+\alpha(t)g\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big],$ (12)
where $\alpha:[0,1]\rightarrow[0,1]$ is an increasing function of the percentage of training progress.
The objectives in (10), (11), and (12) allow us to implement a variety of preference optimization algorithms, while benefiting from a theoretical justification. In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform baselines. In Appendix˜E, we discuss why we chose Bregman divergences rather than $f$-divergences.